# One-Shot Graph Representation Learning Using Hyperdimensional Computing

### v1

@@ 4 Experiments
We proceed to report results of extensive experiments comparing HDGl with against a set of state-of-the-art node labeling methods, including several GNN models, in terms of both accuracy and runtime.
@@ 4.1 Data sets
For our experiments, we use the three of the most commoly used node labeling benchmarks: CORA McCallum et al. (2000), CiteSeer Giles et al. (1998), PubMed Namata et al. (2012) data sets and two data sets from the the Microsoft Academic Graph: Coauthor CS and Coauthor Physics from Shchur et al. (2018). We used the Train/Validation/Test splits described inYang et al. (2016); Kipf and Welling (2017); Veličković et al. (2018) for CORA, CiteSeer and PubMed data sets and that described in Shchur et al. (2018) for Coauthor CS and Coauthor Physics data sets. The data sets are summarized in Table 1.
@@ 4.2 Models
We compare the performance of HDGL with that of several strong baselinesL Multilayer Perceptron (MLP), DeepWalk Perozzi et al. (2014), Label Propagation. We also directly compare HDGL with state-of-the-art models: Graph Convolutional Network (GCN) Kipf and Welling (2017), Graph Attention Network (GAT) Veličković et al. (2018), and lastly, an fast GNN, namely, Simplified Graph Convolutional Network (SGC) Wu et al. (2019) which is essentially a logistic regression model trained using features extracted from $k$-hop neighborhoods of nodes as input.
@@ 4.3 HDGL Model Specifications
Because Bundling of an even number of HD vectors can result in ties that would need to be broken using tie-breaking policy (See Section 2), we opt to randomly sample an odd number, specifically, 11 of the 1-hop neighbors and 21 of the 2-hop neighbors during the construction of HD node representations by HDGL.
In the case of CORA and CiteSeer datasets, we map the node features to $50,000$-dimensional HD vectors via random projections. For the PubMed, Coauthor CS, and Coauthor Physics datasets, we map the node features $20,000$-dimensional HD vectors in a similar fashion. While these choices may seem arbitrary, we note that a choice between $20000$ and $50000$ is typically large enough for the desired properties of HD-computing to hold.
@@ 4.4 Training and Evaluation
Recall that iterative machine learning methods, including GNN typically make use of a validation data set that is distinct from the test data set for deciding when to terminate training, or for tuning the hyper-parameters. Hence, effectively, they make use of both training and validation data sets for building the predictive model and test data for evaluating the trained model’s performance. In contrast, one-shot learning methods such as HDGL have no need for validation data. Hence, to ensure fair comparison between HDGL and other methods, we evaluate their performance on the training data, while letting the individual methods use training and validation data as appropriate for the algorithm in question. Thus, HDGL uses training and validation sets together to build the model, whereas GNN use training data for training the model, validation data for early stopping and hyperparameter tuning. In the case of GNN methods that rely on iterative training, we choose the models with hyperparameters tuned to obtain optimal performance on validation data. For tuning the hyperparameters, we used grid search over a set of 12 configurations. In the case of GCN and GAT, the following hyperparameters were optimized the number of hidden units in the first layer, learning rate, and weight decay. In the case of GCN,networks with 32 and 64 hidden units were explored whereas in the case of GAT, 4 and 8 attention heads, each with 8 hidden units were explored. Both GCN and GAT were explored with weight decays in [1e-3, 1e-4, 5e-4]. In the case of SGC which does not use any hidden layers, only the learning rate and weight decay parameters were tuned in [1e-3, 5e-4, 1e-4, 5e-5, 1e-5, 5e-6]. GCN, GAT, and SGC were tuned for learning rates set to 1e-2 and 1e-3. The early stopping criterion optimized over choices of 25 and 50 epochs, respectively, for all three models. As MLP, DeepWalk, and Label Prop represent well-established baselines, we do not conduct a hyperparameter search. Instead, we choose to utilize publicly released implementations that already incorporate optimal hyperparameters.
All of the models were evaluated on the same test data. In each case, we report measured performance on the test data, averaged over 10 random choices of the relevant parameters: random hyperplanes used to construct HD representations used by HDGL, weight initializations of the deep learning models.
Method Cora Citeseer Pubmed CompSci Physics
MLP 55.1 $\pm$ 0.6 46.5 $\pm$ 0.4 71.4 $\pm$ 0.5 87.9 $\pm$ 0.6 88.2 $\pm$ 1.0
DeepWalk 67.2 $\pm$ 0.8 43.2 $\pm$ 0.6 65.3 $\pm$ 0.5 78.1 $\pm$ 0.9 87.6 $\pm$ 0.9
Label Prop 68.2 $\pm$ 1.1 45.3 $\pm$ 0.8 63.1 $\pm$ 0.5 74.1 $\pm$ 3.9 86.1 $\pm$ 2.1
GCN 81.9 $\pm$ 0.6 70.3 $\pm$ 0.5 79.0 $\pm$ 0.4 91.1 $\pm$ 0.8 92.8 $\pm$ 0.5
GAT 82.8 $\pm$ 0.5 71.0 $\pm$ 0.6 79.1 $\pm$ 0.3 90.5 $\pm$ 0.6 92.5 $\pm$ 0.9
SGC 79.8 $\pm$ 0.0 68.5 $\pm$ 0.1 75.4 $\pm$ 0.0 89.1 $\pm$ 0.1 92.1 $\pm$ 0.1
HDGL 79.5 $\pm$ 0.7 70.7 $\pm$ 0.6 77.8 $\pm$ 1.1 88.9 $\pm$ 0.4 91.0 $\pm$ 0.9
Table 2: Average Test Set Accuracy and Standard Deviation (in Percentage) Over 10 Random Weight Initializations for Deep Learning Models and 10 Randomized Instantiations of HDGL Hyperplanes Across All Datasets
@@ 4.5 Implementation Details
All experiments were conducted on Google Colab Pro with High-memory CPU. To ensure fair comparison, although not ideal for HDGL, we used the Python PyTorch framework Paszke et al. (2019), augmented with the Deep Graph Library (DGL) Wang et al. (2019) for importing datasets and implementing deep learning models. None of our experiments used parallel execution or GPUs.
Figure 2: Comparison of Learning Times (in seconds) Across Datasets for HD-GL with GCN, GAT and SGC with Hyperparameter Tuning with Exploration of Search Space with 12 configurations.
@@ 4.6.1 Predictive Performance
The results of our comparison of HDGL with the other methods are shown in Table 2. We observe that HDGL, an one-shot learning method without the need for computationally expensive iterative training or hyperparameter optimization achieves node prediction accuracy that is competitive with those of state-of-the-art GNN models (GCN, GAT, SGC) across all 5 benchmark data sets, in each case reaching performance that is within 1-2% of the best performing model. HDGL also consistently outperforming basic deep learning baseline methods (MLP, DeepWalk, LabelProp).
4.6.2 Run-times
Fig 2 compares run-time of HDGL (which requires no hyperparameter tuning) with those of GCN, GAT, and SGC (with hyper-parameter tuning) with the run-time of HDGL. We observe that barring one exception, PubMed data where the run-times of SGC and HDGL are close, the run-time of HDGL is substantially less than that of all the state-of-the-art GNN.
The speedup obtained by HDGL over the GNN methods is particularly noticeable for larger graph datasets, such as Coauthor CS and Physics. Figure 3, shows how the methods compare with respect to both predictive accuracy achieved and the run-time used to achieve it. It is apparent that HDGL achieves a substantial speedup by factors of 3.5x to 4 with only minimal (1-2%) drop in performance compared to GNN models.
In practice, deep learning models are often optimized using a far more extensive search over the space of hyperparameters, with corresponding increases in run-time, relative to HDGL, with the use of GPUs further widening the gap between GNN and HDGL in terms of their energy footprints.
Figure 3: Tradeoff of Predictive performance against Run-time for HDGL and GNN on Coauthor CS and Coauthor Physics Datasets. Note that GNN models (GCN, GAT, SGC) performance shown is achieved with hyperparameter tuning whereas HDGL offers one-shot learning without the need for hyperparameter tuning
@@ 5 Conclusion
We have presented a novel simple, one-shot, and hence fast, and efficient approach to semi-supervised algorithm for graph representation learning. The proposed approach constructs low precision, noise-tolerant, similarity preserving randomized hyper-dimensional representations of nodes in graphs. We take advantage of the close connection between the Weisfeiler-Lehman test for graph isomorphism and a family of graph neural networks to design a hyper-dimensional graph learning algorithm. The proposed algorithm uses hyper-dimensional operators such as bundling and binding to aggregate information from the local neighborhood of each node so as to represent nodes using information that is an approximation of that used by the Weisfeiler-Lehman test.
Results of experiments with widely used benchmark data sets show that the proposed algorithm achieves performance on semi-supervised node labeling tasks that is competitive with the state-of-the-art deep learning methods, without the need for computationally expensive optimization of model parameters and hyper-parameters. Consequently, the proposed method offers an especially attractive alternative with energy footprint that is substantially smaller than that of deep learning based modern graph embedding methods.
Because the model supports one-shot learning of node embeddings, it lends itself to use in settings that call for continual learning algorithms De Lange et al. (2021); Wickramasinghe et al. (2023) wherein data samples, or classes, are presented on an ongoing basis, allowing the model to be continually updated, without the need for retraining on old data. Additional directions for further research include variants of the proposed algorithm that can learn graph representations that invariant or equivariant with respect to a specified set of operations on graphs Satorras et al. (2021); Keriven and Peyré (2019).

### v2

@@ 5 Experiments
We proceed to report results of extensive experiments comparing HDGL with a set of state-of-the-art node classification methods and link prediction methods, including several GNN models, in terms of both predictive performance and runtime.
@@ 5.1 Data sets
Since HDGL is designed in such a way that nodes with similar features and neighborhoods will have similar representations, we exclusively concentrate on homophilous graphs in this study, deferring the exploration of more complex cases involving heterophily graphs to future research.
For our experiments, we utilize commonly used node labeling homophilous graph benchmarks: the CORA graph dataset [61], sourced from McCallum et al. [32]; the CiteSeer graph dataset [61], sourced from Giles et al. [16]; and the PubMed dataset [36]. Additionally, we include three more datasets from the Microsoft Academic Graph: Coauthor CS and Coauthor Physics from Shchur et al. [45], and the BlogCatalog Graph dataset from Huang et al. [20]. Additionally, we introduce a new version of the DBLP dataset, built upon the works of Pan et al. [39]. We remove nodes with no neighbors in the graph obtained from Pan et al. [39]. The original node features from Pan et al. [39] consist of manuscript titles. We utilize BERT [11] to obtain embeddings for the titles and employ them as node features. The data sets and their statistics are given in Table 1.
We adhere to the Train/Validation/Test splits described in Yang et al. [61], Kipf and Welling [26], and Veličković et al. [53] for the CORA, CiteSeer, and PubMed datasets, respectively. For the Coauthor CS, Coauthor Physics, BlogCatalog, and DBLP datasets, we employ the data splitting strategy outlined in Shchur et al. [45].
We test link prediction models on CORA, CiteSeer, and Pubmed datasets. These models are trained on incomplete data, retaining node features but omitting some edges. Validation and test sets include excluded edges and randomly selected unconnected node pairs. Evaluation assesses the models’ ability to classify both edges and non-edges. We follow splitting strategy from Kipf and Welling [25] work i.e Validation and test sets each contain 5% and 10% of citation links.
@@ 5.2 Models
For semi-supervised node classifcation task, we compare the performance of HDGL with that of several strong baselines Logistic Regression (LogReg), DeepWalk [42] and Label Propagation [64]. We also directly compare HDGL with state-of-the-art models: Graph Convolutional Network (GCN) [26], Graph Attention Network (GAT) [53], and lastly, a fast GNN model, namely, Simplified Graph Convolutional Network (SGC) [57] which is essentially a logistic regression model trained using features extracted from $k$-hop neighborhoods of nodes as input. We also assess our model against RelHD [24], an existing HD algorithm used for node classification.
For the transductive link prediction task, we evaluate our model against standard link prediction baselines:- DeepWalk [42] and Spectral clustering (SC) [48]. We also compare our model against graph autoencoder models:- Graph Autoencoder (GAE) and Variational Graph Autoencoder (VGAE) from Kipf and Welling [25].
@@ 5.3 HDGL Model Specifications
Because Bundling of an even number of HD vectors can result in ties that would need to be broken using tie-breaking policy (See Section 3.2), we opt to randomly sample an odd number, specifically, 11 of the 1-hop neighbors and 21 of the 2-hop neighbors during the construction of HD node representations by HDGL for datasets except BlogCatalog. For BlogCatalog we sample 21 of the 1-hop neighbors and 51 of the 2-hop neighbors since Blogcatalog graph is denser than the other datasets.
For node classification in CORA, CiteSeer, and BlogCatalog datasets, we map node features to $50,000$-dimensional HD vectors through random projections. Similarly, for the PubMed, Coauthor CS, Coauthor Physics, and DBLP datasets, we map node features to $20,000$-dimensional HD vectors using the same method. While these choices may seem arbitrary, it’s crucial to note that selecting dimensions between $20,000$ and $50,000$ typically suffices to maintain the desired properties of HD-computing. This assertion finds extensive support in the research from Kanerva [22, 23].
Table 2: Average Test Set Accuracy and Standard Deviation (in Percentage) Over 10 Random Weight Initializations for Deep Learning Models and 10 Randomized Instantiations of RelHD and HDGL Across Datasets (* denotes datasets with $\{0,1\}$ features. In these datasets, RelHD performs comparably to HDGL)
Method Cora Citeseer Pubmed BlogCat CompSci Physics DBLP
LogReg 52.2 $\pm$ 0.5 46.3 $\pm$ 0.4 67.9 $\pm$ 0.5 65.3 $\pm$ 1.0 86.8 $\pm$ 0.6 88.2 $\pm$ 0.9 56.2 $\pm$ 3.0
DeepWalk 65.3 $\pm$ 1.2 46.2 $\pm$ 0.9 69.3 $\pm$ 0.9 58.6 $\pm$ 1.3 78.1 $\pm$ 0.9 86.4 $\pm$ 0.9 66.1 $\pm$ 2.2
Label Prop 70.9 $\pm$ 0.0 47.7 $\pm$ 0.0 71.3 $\pm$ 0.0 47.8 $\pm$ 3.8 76.6 $\pm$ 0.9 85.6 $\pm$ 1.1 66.4 $\pm$ 1.6
GCN 81.9 $\pm$ 0.6 70.3 $\pm$ 0.5 79.0 $\pm$ 0.4 69.5 $\pm$ 0.4 91.0 $\pm$ 0.7 92.3 $\pm$ 0.6 72.1 $\pm$ 0.8
GAT 82.8 $\pm$ 0.5 72.0 $\pm$ 0.6 79.1 $\pm$ 0.3 65.3 $\pm$ 1.2 90.9 $\pm$ 0.6 92.5 $\pm$ 0.9 75.4 $\pm$ 1.0
SGC 79.8 $\pm$ 0.5 68.5 $\pm$ 0.4 75.4 $\pm$ 0.5 70.8 $\pm$ 0.3 89.1 $\pm$ 0.6 92.1 $\pm$ 0.7 76.3 $\pm$ 0.9
RelHD 78.9 $\pm$ 0.5 69.3 $\pm$ 0.4 69.1 $\pm$ 0.4 60.9 $\pm$ 1.2 85.5 $\pm$ 0.8 85.3 $\pm$ 0.9 44.9 $\pm$ 0.2
HDGL 79.5 $\pm$ 0.7 70.0 $\pm$ 0.6 76.8 $\pm$ 1.1 69.3 $\pm$ 0.5 88.9 $\pm$ 0.4 91.0 $\pm$ 0.9 68.7 $\pm$ 1.2
@@ 5.4 Training and Evaluation
Recall that iterative machine learning methods, including GNN typically make use of a validation data set that is distinct from the test data set for deciding when to terminate training, or for tuning the hyper-parameters. Hence, effectively, they make use of both training and validation data sets for building the predictive model and test data for evaluating the trained model’s performance. In contrast, learning methods such as HDGL which invole no iterative training do not need validation data. Hence, to ensure fair comparison between HDGL and other methods, we evaluate their performance on the training data, while letting the individual methods use training and validation data as appropriate for the algorithm in question. Thus, HDGL uses training and validation sets together for learning, whereas GNN use training data for training the model, validation data for early stopping and hyperparameter tuning.
In the case of GNN methods that rely on iterative training, we choose the models with hyperparameters tuned to obtain optimal performance on validation data. For tuning the hyperparameters, we used grid search over a set of 12 configurations. In the case of GCN and GAT, the following hyperparameters were optimized the number of hidden units in the first layer, learning rate, and weight decay. In the case of GCN,networks with 32 and 64 hidden units were explored whereas in the case of GAT, 4 and 8 attention heads, each with 8 hidden units were explored. Both GCN and GAT were explored with weight decays in [1e-3, 1e-4, 5e-4]. In the case of SGC which does not use any hidden layers, only the learning rate and weight decay parameters were tuned in [1e-3, 5e-4, 1e-4, 5e-5, 1e-5, 5e-6]. GCN, GAT, and SGC were tuned for learning rates set to 1e-2 and 1e-3. The early stopping criterion optimized over choices of 25 and 50 epochs, respectively, for all three GNN models.
As LogReg, DeepWalk, and Label Prop represent well-established baselines, we do not conduct a hyperparameter search. For Label Propagation, we use 20 iterations of propagation (except for blogcatalog dataset where we use 10) and set the alpha factor  [^1] Alpha factor in Label Propagation determines the balance between existing and propagated labels. Higher alpha prioritizes existing labels; lower alpha favors propagated ones. to 0.5. For DeepWalk, we utilize the DGL library, setting the embedding dimension to 32 and the number of walks to 20, with a batch size of 64. We sample such walks for all nodes in the graph and train DeepWalk model for 10 epochs and then employ logistic regression for classification, incorporating early stopping based on the validation set.
Following Kang et al. [24], the RelHD algorithm was employed with 10,000 dimensional random basis HD vectors, as the performance of RelHD plateaus around this dimensionality.
We report measured performance on the test data, averaged across 10 random selections of pertinent parameters. These parameters include the random hyperplanes employed for constructing high-dimensional representations used by HDGL, random basis vector initializations for RelHD, weight initializations of the deep learning models, and data splits for datasets such as Coauthor Physics, Coauthor CS, BlogCatalog, and DBLP. For Cora, Citeseer, and Pubmed, we use predefined splits given in Yang et al. [61].
For the link prediction task, we rely on the results directly obtained from the paper for GVAE from Kipf and Welling [25].
@@ 5.5 Implementation Details
All experiments were conducted on Google Colab Pro with High-memory CPU. To ensure fair comparison, although not ideal for HDGL, we used the Python PyTorch framework [40], complemented by the Deep Graph Library (DGL) [54] for importing datasets and implementing deep learning models. None of our experiments used parallel execution or GPUs. We have also developed a PyTorch version of RelHD, as the codebase was not publicly available and experiments detailed in Kang et al. [24] were reliant on hardware architecture implementations. We undertake this effort to compare with our model and use RelHD as a baseline in this study.
@@ 5.6 Results
Node Classification Performance. The results of our comparison of HDGL with the other methods are shown in Table 2. We observe that HDGL without the need for computationally expensive iterative training or hyperparameter optimization achieves node prediction accuracy that is competitive with those of state-of-the-art GNN models (GCN, GAT, SGC) across most of the benchmark data sets, in each case reaching performance that is within 1-2% of the best performing model. HDGL also consistently outperforming basic deep learning baseline methods (LogReg, DeepWalk, LabelProp). The notable advantage of HDGL over RelHD lies in its ability to accommodate features beyond the binary $\{0,1\}$ range, as evident from the results presented in Table 2, where HDGL significantly outperforms. This is particularly notable with datasets lacking $\{0,1\}$ features. However, when dealing with datasets featuring $\{0,1\}$ attributes like CORA and Citeseer, the performance of HDGL and RelHD appears nearly identical.
To highlight the advantage of HDGL more prominently, we conducted experiments for node classification on the CORA and CiteSeer datasets where we reduced the node features to 100 dimensions using PCA [41], thereby trying to enforce continuous node features. As depicted in Figure 2, the impact is evident through a noticeable drop in performance for RelHD. Conversely, for other methods including HDGL, the performance remains relatively unchanged.
Figure 2: Impact of performing PCA with a dimensionality of 100 on features for node classification accuracy.
Figure 3: Comparison of Learning Times (in seconds) i.e Across Datasets for HDGL and RelHD with GCN, GAT and SGC with Hyperparameter Tuning with Exploration of Search Space with 12 configurations. Runtime for HDGL and RelHD exclusively includes only time Elapsed for learning label representations using train/validation data.
Table 3: Link prediction results averaged over 10 Randomized Instantiations of HDGL and datasplits. Results for other models are from Kipf and Welling [25] paper.
Methods Cora Citeseer PubMed
AUC AP AUC AP AUC AP
DeepWalk 0.831 0.850 0.805 0.836 0.844 0.841
SC 0.846 0.885 0.791 0.826 0.849 0.888
GAE 0.910 0.920 0.895 0.899 0.964 0.965
VGAE 0.914 0.926 0.908 0.920 0.944 0.947
HDGL 0.849 0.880 0.768 0.842 0.853 0.884
Node Classification Run-times. Figure 3 compares the run-time for the node classification task between HDGL and RelHD (both of which requires no hyperparameter tuning) with those of GCN, GAT, and SGC (with hyperparameter tuning), alongside the run-time of HDGL. We observe that, the run-time of HDGL is substantially less than that of all the state-of-the-art GNNs. While in most cases, RelHD exhibits slightly or significantly worse run times compared to HDGL, it’s crucial to note that HDGL consistently delivers superior performance across various datasets when compared to RelHD.
Table 4: Class Incremental Node Labeling Results. No hyperparameter grid search is used for GCN and SGC algorithms. Wall-time includes both training and inference of the model. HDGL does not require retraining from $t=2$.
Model BlogCatalog Coauthor Physics
t = 1 t = 2 t = 3 t = 4 t = 5 t=1 t=2 t=3 t=4
GCN Accuracy 94.94 85.81 75.16 74.82 69.61 98.95 97.33 95.57 92.58
Runtime 51.2s 18.7s 16.5s 33.9s 23.9s 65.9s 62.3s 65.8s 51.2s
SGC Accuracy 94.56 82.31 73.23 72.72 69.81 98.94 97.39 95.36 91.81
Runtime 6.3s 3.72s 3.43s 3.73s 3.6s 9.1s 8.5s 8.8s 9.7s
HDGL Accuracy 89.11 88.49 75.56 71.73 69.91 98.56 94.87 92.53 90.8
Runtime 20.2s 0.3s 0.4s 0.4s 0.6s 38.3s 1.5s 1.7s 1.7s
It is crucial to note that while we have used grid search on 12 potential setups for the deep learning model hyperparameter optimization of deep learning models typically explores a much larger space. This expanded search naturally comes with increased runtime, especially when GPUs are utilized, further widening the energy efficiency gap between Graph Neural Networks (GNN) and traditional deep learning models like HDGL.
Link Prediction Performance. Even though HDGL representations are primarily tailored for transductive node classification, their performance for link prediction is comparable to that of DeepWalk and Spectral clustering; as seen from Table 3. However, they do not reach the same level of effectiveness as GAE and VGAE, which leverage GNN methods. Despite lacking iterative training like deep learning methods, HDGL still manages to achieve respectable results in link prediction. Moreover, unlike deep learning approaches that typically require distinct models for link and node classification, HDGL tackles both tasks within a unified framework. We also perform link prediction experiments by varying the dimensionality of the HDGL model. This exploration arises from observing a slight yet discernible variation in performance for link prediction tasks, unlike the more consistent outcomes observed in node classification experiments. As mentioned earlier, typically, dimensions ranging from 20,000 to 50,000 are sufficient for HD properties to hold. However, higher dimensionality generally leads to better performance. Nevertheless, beyond a certain threshold, the performance tends to plateau, as illustrated in Figure 4.
Figure 4: Link Prediction Performance of HDGL under Various dimensionality configurations.
HDGL in the Class-Incremental Learning Setting Consider a scenario with a dynamic graph where nodes and edges remain static, but the new node labels appear over time i.e initially, only a subset of nodes is labeled, belonging to two classes. Over time, new nodes receive labels; thus, introducing new classes. This is a node labeling instance of the class-incremental learning problem [52, 3, 51]. In this setting, deep learning methods typically require either retraining on the data for previously learned classes along with the data for the new classes to avoid catastrophic forgetting of the previously learned classes [51]. This retraining incurs significant additional runtime and computation.
In contrast to deep neural networks, which necessitate retraining, Hyperdimensional Graph Learner (HDGL) operates without an iterative training phase. Instead, it constructs label hypervectors using predefined latent node representations. This approach eliminates the need for retraining, resulting in substantial savings in both runtime and computation costs.
To simulate incremental node classification on a fully labeled graph, we use BlogCatalog and Coauthor Physics Dataset. We first split the nodes into train/val/test as in Shchur et al. [45]. We mask the train/val/test such that at time $t=1$ only the labels $\{1,2\}$ are available; and at $t=2$ only the labels $\{1,2,3\}$ are only available; and so on such that at the final time step, all of the labels $\{1,\cdots,L\}$ are available.
To simplify matters for GNN models, we assume that the models are provided the list of $L$ possible labels although only a subset have been encountered in the training data. This allows the last layer of the GNN to have $L$ outputs. Also, we forego hyperpameter search for the GNN models to minimize the computational overhead.
In comparing HDGL against GCN and SGC in our experimental setup, we opted not to include the GAT model due to its longer runtime, as depicted in Figure 3. As seen from Table 4, we see that HDGL incurs a significant computational overhead only at timestep 1, necessitating the computation of node representations followed by inference. However, for subsequent timesteps , HDGL leverages previously computed node representations, requiring only inference. In contrast, deep learning models like GCN and SGC mandate retraining for each timestep to accommodate new labels, resulting in comparatively slower runtimes.
@@ 6 Conclusion
We introduced HDGL, a novel and efficient algorithm for transductive learning on graphs using hyperdimensional representations. HDGL leverages the injectivity property of node representations of a family of Graph Neural Networks (GNNs) to map node features to the HD space and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node. The resulting latent node representations support both node classification and link prediction tasks, unlike typical deep learning methods, which often require separate models for these tasks. HDGL obviates the need for expensive iterative training of the sort required by deep learning methods. HDGL is also well-suited for class incremental learning where the model has to learn to effectively discriminate between a growing number of classes.
Our experiments with widely used benchmark datasets show that on the node classification task, HDGL is competitive with the state-of-the-art GNN methods with respect to accuracy, at substantially reduced computational cost. Our experiments also show that the HD representation constructed by HDGL supports link prediction at accuracies comparable to that of DeepWalk and related methods. We conclude that HDGL offers a computationally efficient alternative to graph neural networks for node classification, especially in settings that call for class-incremental learning or in applications that demand high accuracy models at significantly lower computational cost and learning time than possible with the state-of-the-art graph neural network methods.